OpenStudy (pumpkin1102):

The work of a student to solve a set of equations is shown below: Equation 1: m = 6 + n Equation 2: 2m = 8 + 8n Step 1: −2(m) = −2(6 + n) 2m = 8 + 8n Step 2: −2m = −12 − 2n 2m = 8 + 8n Step 3: −12 − 2n = 8 + 8n Step 4: −20 = 10n Step 5: n = −2 In which step did the student first make an error? Step 2 Step 3 Step 4 Step 5

OpenStudy (welshfella):

Hint there is a mistake in adding the 2 equations

OpenStudy (welshfella):

sorry gotta go now

OpenStudy (pumpkin1102):

either a or b

OpenStudy (pumpkin1102):

I think

OpenStudy (pumpkin1102):

plz help

OpenStudy (ciarán95):

It looks to be in Step 3 I think. The student seems to let the two equations (from Step 2) equal one another rather than adding them together. In doing so, they would have to be assuming that -2m = 2m (i.e. m = 0), which cannot be assumed here when solving these equations. -2m = -12 - 2n 2m = 8 + 8n If we add these two equations together, we should get: 0 = -4 + 6n => 6n = 4 => n = 6/4

OpenStudy (pumpkin1102):

thx